Session: Mar 2010 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Thu Mar 18 13:50:36 2010) Success:

Shanghai (Bo Xia at Thu Mar 18 13:52:38 2010) Success:

Onsala (Michael Lindqvist at Thu Mar 18 13:58:40 2010) Minor failures:
Very foggy. Problem with MPI-0606. Unstable green lights on the Mark5, flickering while recording. The same pack was used in eb039fon, see comments. Will not use that for 1 Gbps recording. But is a 6 TB pack with about 70% free :-(.

Yebes (Pablo de Vicente at Thu Mar 18 14:21:09 2010) Success:
No known problems

Effelsberg (Uwe Bach at Thu Mar 18 14:27:07 2010) Success:

Westerbork (Geert Kuper at Thu Mar 18 19:40:34 2010) Success:
12 tescopes in array. No known problems.

Torun (K. Borkowski at Fri Mar 19 06:36:12 2010) Minor failures:
Succeeded to repair the C-band receiver virtually the last minute. The sensitivity seems as in the past. Unfortunately there were no noise diode signals! As there were nice fringes we decided to continue this experiment and get to repair the diode afterwards.

Noto (P. Cassaro at Fri Mar 26 09:33:55 2010) Success:

Jodrell_Bank (alastair gunn at Mon Mar 29 12:22:21 2010) Success:
No known problems.

Cambridge (alastair gunn at Mon Mar 29 12:24:07 2010) Success:
Kn on BBCs 1 and 2. No known problems.